# Bilevel Learning for Bilevel Planning

### v2

@@ Abstract
A robot that learns from demonstrations should not just imitate what it sees—it should understand the high-level concepts that are being demonstrated and generalize them to new tasks. Bilevel planning is a hierarchical model-based approach where predicates (relational state abstractions) can be leveraged to achieve compositional generalization. However, previous bilevel planning approaches depend on predicates that are either hand-engineered or restricted to very simple forms, limiting their scalability to sophisticated, high-dimensional state spaces. To address this limitation, we present IVNTR, the first bilevel planning approach capable of learning neural predicates directly from demonstrations. Our key innovation is a neuro-symbolic bilevel learning framework that mirrors the structure of bilevel planning. In IVNTR, symbolic learning of the predicate “effects” and neural learning of the predicate “functions” alternate, with each providing guidance for the other. We evaluate IVNTR in six diverse robot planning domains, demonstrating its effectiveness in abstracting various continuous and high-dimensional states. While most existing approaches struggle to generalize (with $<35\%$ success rate), our IVNTR achieves an average of $77\%$ success rate on unseen tasks. Additionally, we showcase IVNTR on a mobile manipulator, where it learns to perform real-world mobile manipulation tasks and generalizes to unseen test scenarios that feature new objects, new states, and longer task horizons. Our findings underscore the promise of learning and planning with abstractions as a path towards high-level generalization. Work was partly done during internship at Centaur AI Institute. Correspondence to {bowenli2,basti}@andrew.cmu.edu.
@@ I Introduction
Imitation learning has made significant recent strides [1, 2, 3, 4, 5], but generalization remains an open challenge, especially when new tasks require recomposing high-level concepts that are only implicit in the training data [6, 7]. In Figure LABEL:fig:teaser, a robot has seen demonstrations of stepping onto a platform to grasp an object, and other demonstrations of dumping the object into a container. Now, faced with a new task where the container is also elevated, the robot should first move the two platforms appropriately, step onto one of them to grasp the object, and finally step on the other to dump the object. Note that the platform arrangements must be completed before grasping the target, since the robot cannot move platforms with its hand full. This kind of learning and reasoning requires compositional generalization (new objects); sequential generalization (new and longer action sequences); and long-horizon planning with continuous state and action spaces with sparse feedback (goals). In sum, the robot should not just imitate demonstrations, but also understand and leverage the high-level concepts within the low-level states that are being demonstrated.
@@
One promising direction to address these challenges is to learn and plan with abstractions [8, 9, 10, 11, 12, 13]. In this work, we continue a line of recent inquiry on learning abstractions for bilevel planning [14, 15, 16, 17, 18, 19, 20]. In bilevel planning, continuous low-level states are mapped into a symbolic relational state space defined by predicates such as Viewable(?robot,?target) or On(?robot,?platform). Planning proceeds jointly in the symbolic high-level space and the continuous low-level space. The key idea is that this hybrid planning can be more efficient and effective than reasoning solely in the low-level space.
@@
The performance of bilevel planning depends substantially on the predicates used to define the abstract state space [16]. To avoid the need for a human engineer to manually define predicates for every new domain, recent work has considered learning predicates from data [21, 22, 16, 23, 24, 20]. Broadly, three approaches have emerged. The most direct one relies on human feedback (labels or guidance) during the predicate learning process [23, 25, 24], which is labor-intensive and does not guarantee useful abstractions for planning [16]. The second approach invents predicates with surrogate objectives that are easy to optimize, e.g., reconstruction loss [26, 27, 28] or bisimulation [12, 29]. While these methods simplify learning, they complicate planning due to the mismatch between the surrogate objectives and the actual planning goals [16]. The third approach directly invents predicates for efficient planning, making planning “easy” but learning “hard,” as objectives like total-planning-time and expected-planning-success are difficult to optimize [16]. To address this, previous works have used program synthesis with classical grammars [16] and foundation model-based techniques [20, 30]. However, in both cases, the predicates are invented from programmatic and pre-defined functions, which are limited in flexibility and scalability.
@@
Our main contribution is bIleVel learNing from TRansitions (IVNTR), the first approach capable of learning neural predicates that are optimized for efficient and effective bilevel planning. Since directly incorporating the planning objective into network training is challenging, our IVNTR instead constructs a candidate neural predicate pool, which is later subselected [16]. The key insight behind our approach is to center learning around the effects of predicates, which provide two major benefits: (1) they enable the derivation of supervision labels for transition pairs, yielding a well-structured learning objective for training the neural network; and (2) the inherent sparsity of predicate effects, combined with neural learning signals, facilitates efficient symbolic learning of their structure. To this end, IVNTR presents an innovative bilevel learning framework, inspired by the structure of bilevel planning itself. Similar to the alternation between high-level symbolic search and low-level neural sampling in bilevel planning, IVNTR interleaves symbolic effect learning and neural function learning in an iterative process. In each iteration, the symbolic learning proposes a candidate predicate effect across different actions, which provides labels for neural learning on transition pairs. Once the neural classifier converges, its validation loss guides the symbolic learning to propose the next candidate that could minimize the loss in the new iteration. This iterative bilevel learning ultimately yields a compact set of neural predicates, which are then selected to optimize the planning objective [16]. The final set of invented predicates seamlessly integrates into operator and sampler learning frameworks [31, 16], ultimately forming a fully functional bilevel planner.
@@
To evaluate the effectiveness of IVNTR, we conduct extensive experiments across six diverse robot planning domains. These domains feature a wide range of low-level state representations, from SE(2) and SE(3) poses to high-dimensional point clouds. Furthermore, as shown in Figure LABEL:fig:teaser, by leveraging relational predicates and AI planning, IVNTR zero-shot generalizes to tasks with unseen entity compositions. Finally, we deploy IVNTR on a quadruped mobile manipulator (Boston Dynamics Spot) for two long-horizon mobile manipulation tasks. The learned predicates successfully abstract complex continuous states into representations compatible with the task planner, while also providing actionable guidance for the motion planner. We believe IVNTR represents a pivotal step towards learning high-level abstractions from sophisticated low-level states.
@@ II Problem Formulation
A solution to a task is a plan $\pi=[\underline{\mathtt{C}}_{1},\underline{\mathtt{C}}_{2},\cdots,\underline{\mathtt{C}}_{H}]$, that is, a sequence of $H$ ground actions such that successive application of the transition model $\mathbf{x}_{i}=f\left(\mathbf{x}_{i-1},\underline{\mathtt{C}}_{i}\right)$ on each $\underline{\mathtt{C}}_{i}\in\pi$, starting from $\mathbf{x}_{0}$, results in a final state $\mathbf{x}_{H}$ where $g$ holds. For instance, the plan depicted in Figure LABEL:fig:teaser right bottom finally leads to the state where $\texttt{In}(\texttt{t}_{1},\texttt{t}_{2})$ holds. In sum, to generate the plan $\pi$ for a task, in each state, the robot needs to predict: (1) the action class $\mathtt{C}\in\mathcal{C}$, (2) the objects as the discrete parameters, and (3) the continuous parameter $\omega\in\Omega$.
During training, the robot has access to an offline demonstration dataset $\mathcal{D}=\{({T}_{i},\pi_{i})\}_{i=1}^{B}$, which consists of $B$ task and solution pairs sampled from the task distribution $\mathcal{T}^{\mathrm{train}}$. Note that since the transition function $f$ is known and deterministic, we can also recover the intermediate states from $\mathbf{x}_{0}$. During test time, the robot is required to solve held-out tasks sampled from a different test distribution $\mathcal{T}^{\mathrm{test}}$. In practice, for the sake of evaluating generalization, test tasks typically contain new and more objects than training tasks. For example, as shown in Figure 1, all training tasks only have $1$ platform, but during test, there are $2$ platforms. The difference in object compositions could result in different lengths of plans with different action order, requiring the method to generalize by understanding the implicit concepts present in the training demonstrations.
@@ III Bilevel Planning
Figure 2: (a) Overview of IVNTR during training. Given transition pairs in the continuous space, IVNTR invents neural predicates group by group, resulting in the candidate set. A subset that minimize the planning objective $J(\cdot)$ is selected from the candidates, which serves as the finial $\Psi_{\mathrm{dyn}}$. With the complete predicate set, sampler and operator learning can be achieved. (b) Bilevel planning with operators and samplers during test. The compositional ground predicates provides search configurations for task planner and sampling guidance for motion planner.
@@
The operator for a parametrized controller $\mathtt{C}$ is a tuple, $\mathtt{Op}^{\mathtt{C}}=\langle\mathtt{Var},\mathtt{Pre},\mathtt{Eff}^{+},\mathtt{Eff}^{-}\rangle$, where $\mathtt{Var}$ is a tuple of object placeholders matching the type signature of $\mathtt{C}$, and $\mathtt{Pre},\,\mathtt{Eff}^{+},\,\mathtt{Eff}^{-}\subseteq\Psi$, respectively preconditions, add effects, and delete effects, are each a set of lifted predicates defined with variables in $\mathtt{Var}$. $\Psi$ is an oracle predicate set.
@@
Given a task $T=\langle\mathcal{O},\mathbf{x}_{0},g\rangle$, bilevel planning (Figure 2b) starts by using predicates to generate an abstract state consisting of all ground predicates with objects $\mathcal{O}$ whose classifiers evaluate to True in $\mathbf{x}_{0}$. The initial ground predicates, together with the operator set and goal $g$, can then be fed into an AI planner [33] to generate a plan skeleton, $\bar{\pi}$ with partially grounded actions with unspecified continuous parameters, $\underline{\bar{\mathtt{C}}}$. To refine the actions in this skeleton $\underline{\bar{\mathtt{C}}}\in\bar{\pi}$ into fully grounded $\underline{{\mathtt{C}}}$ with the continuous parameters $\omega$, bilevel planning leverages samplers.
@@
Note that unlike the deterministic operators, samplers are usually stochastic and may fail in certain steps. Thus, bilevel planning alternates between the task planner and motion planner and uses the predicate functions $\theta_{\Psi}$ as “guidance” in each step to compensate for the potential sampling failure.
Assuming we have the complete predicate set and their classifiers, previous work has studied the problem of learning operators [31] and samplers [18, 15] from the demonstration dataset $\mathcal{D}^{\mathrm{demo}}$. Since the predicates, learned operators, and samplers are relational, they can be generally applied to held out test tasks sampled from $\mathcal{T}^{\mathrm{test}}$. However, with an insufficient predicate set—for example, with only $\Psi_{\mathrm{G}}$ and $\Psi_{\mathrm{sta}}$—bilevel planning can be intractably slow [16]. We next introduce the IVNTR framework that closes this gap by automatically inventing dynamic predicates for efficient bilevel planning.
@@ IV Methodology
The problem of inventing dynamic predicates $\Psi_{\mathrm{dyn}}$ can be decomposed into symbolic learning—how many predicates should be invented, and with what type signatures—and classifier learning, determining $\theta_{\psi}$ for each invented predicate $\psi\in\Psi_{\mathrm{dyn}}$. Previous approaches [20, 16] address these problems via a “define-then-select” two-stage pipeline. In the first stage, for each possible typed predicate candidate $\hat{\psi}$, its function is pre-defined via program synthesis [16] or pre-trained foundation models [20]. These candidates form a large predicate pool $\hat{\Psi}_{\mathrm{dyn}}$. In the second stage, to subselect predicates from the pool, each candidate predicate set $\tilde{\Psi}_{\mathrm{dyn}}\subseteq\hat{\Psi}_{\mathrm{dyn}}$ is scored with a function $J(\tilde{\Psi}_{\mathrm{dyn}})$ that measures both planning efficiency and effectiveness. A key limitation of this “define-then-select” pipeline is that the classification functions within $\hat{\Psi}_{\mathrm{dyn}}$ are restricted to a relatively simple set. Scaling to more general classifiers, e.g., neural networks, is nontrivial, since $J(\tilde{\Psi}_{\mathrm{dyn}})$ is highly non-differentiable. To address this, we propose IVNTR, a “learn-then-select” approach that leverages bilevel learning.
As depicted in Figure 2 (a), given the domain types $\Lambda$, IVNTR enumerates all possible typed variable compositions (with maximum input arity). Since predicates with different typed variables take different input object features, IVNTR invents them group by group independently. For the invention of each group, IVNTR draws inspiration from bilevel planning itself, where planning alternates between the symbolic level and the low level. Similarly, IVNTR interleaves symbolic learning and neural learning, with each providing guidance for the other. Specifically, symbolic learning proposes effect vectors that represent the add and delete effects for candidate predicates across all operators. Neural learning uses these effect vectors to provide supervision for classifier learning. The validation loss in neural learning feeds back into symbolic learning, and the process repeats. In this section, as shown in Figure 3, we describe these steps in detail via the exemplar predicate group $\psi\in\Psi^{\mathtt{Var}}$, with the typed variables $\mathtt{Var}=\mathtt{(?r,?t)}$.

### v3

@@ Abstract
A robot that learns from demonstrations should not just imitate what it sees—it should understand the high-level concepts that are being demonstrated and generalize them to new tasks. Bilevel planning is a hierarchical model-based approach where predicates (relational state abstractions) can be leveraged to achieve compositional generalization. However, previous bilevel planning approaches depend on predicates that are either hand-engineered or restricted to very simple forms, limiting their scalability to sophisticated, high-dimensional state spaces. To address this limitation, we present IVNTR, the first bilevel planning approach capable of learning neural predicates directly from demonstrations. Our key innovation is a neuro-symbolic bilevel learning framework that mirrors the structure of bilevel planning. In IVNTR, symbolic learning of the predicate “effects” and neural learning of the predicate “classifiers” alternate, with each providing guidance for the other. We evaluate IVNTR in six diverse robot planning domains, demonstrating its effectiveness in abstracting various continuous and high-dimensional states. While most existing approaches struggle to generalize (with $<35\%$ success rate), our IVNTR achieves an average success rate of $77\%$ on unseen tasks. Additionally, we showcase IVNTR on a mobile manipulator, where it learns to perform real-world mobile manipulation tasks and generalizes to unseen test scenarios that feature new objects, new states, and longer task horizons. Our findings underscore the promise of learning and planning with abstractions as a path towards high-level generalization. Project website: https://jaraxxus-me.github.io/IVNTR/. Work was partly done during internship at Centaur AI Institute. Correspondence to {bowenli2,basti}@andrew.cmu.edu.
@@ I Introduction
Imitation learning has made significant recent strides [1, 2, 3, 4, 5], but generalization remains an open challenge, especially when new tasks require recomposing high-level concepts that are only implicit in the training data [6, 7]. In Figure LABEL:fig:teaser, a robot has seen demonstrations of stepping onto a platform to grasp an object, and other demonstrations of dropping the object into a container. Now, faced with a new task where the container is also elevated, the robot should first move the two platforms appropriately, step onto one of them to grasp the object, and finally step onto the other to drop the object. Note that the platform arrangements must be completed before grasping the target, since the robot cannot move platforms with its hand full. This kind of learning and reasoning requires compositional generalization (new objects); sequential generalization (new and longer action sequences); and long-horizon planning with continuous state and action spaces with sparse feedback (goals). In sum, the robot should not just imitate demonstrations, but also understand and leverage the high-level concepts within the low-level states that are being demonstrated.
@@
One promising direction to address these challenges is to learn and plan with abstractions [8, 9, 10, 11, 12, 13, 14]. In this work, we continue a line of recent inquiry on learning abstractions for bilevel planning [15, 16, 17, 18, 19, 20, 21]. In bilevel planning, continuous low-level states are mapped into a symbolic relational state space defined by predicates such as Viewable(?robot,?target) or On(?robot,?platform). Planning proceeds jointly in the symbolic high-level space and the continuous low-level space. The key idea is that this hybrid planning can be more efficient and effective than reasoning solely in the low-level space.
@@
The performance of bilevel planning depends substantially on the predicates used to define the abstract state space [17]. To avoid the need for a human engineer to manually define predicates for every new domain, recent work has considered learning predicates from data [22, 23, 17, 24, 25, 21]. Broadly, three approaches have emerged. The most direct one relies on human feedback (labels or guidance) during the predicate learning process [24, 26, 25], which is labor-intensive and does not guarantee useful abstractions for planning [17]. The second approach invents predicates with surrogate objectives that are easy to optimize, e.g., reconstruction loss [27, 28, 29] or bisimulation [12, 30]. While these methods simplify learning, they complicate planning due to the mismatch between the surrogate objectives and the actual planning goals [17]. The third approach directly invents predicates for efficient planning, making planning “easy” but learning “hard,” as objectives like total-planning-time and expected-planning-success are difficult to optimize [17]. To address this, previous works have used program synthesis with classical grammars [17] and foundation model-based techniques [21, 31]. However, in both cases, the predicates are invented from programmatic and pre-defined classifiers, which are limited in flexibility and scalability.
@@
Our main contribution is bIleVel learNing from TRansitions (IVNTR), the first approach capable of learning neural predicates that are optimized for efficient and effective bilevel planning. Since directly incorporating the planning objective into network training is challenging, our IVNTR instead constructs a candidate neural predicate pool, which is later subselected [17]. The key insight behind our approach is to center learning around the effects of predicates, which provide two major benefits: (1) they enable the derivation of supervision labels for transition pairs, yielding a well-structured learning objective for training the neural network; and (2) the inherent sparsity of predicate effects, combined with neural learning signals, facilitates efficient symbolic learning of their structure. To this end, IVNTR presents a novel bilevel learning framework, inspired by the structure of bilevel planning itself. Similar to the alternation between high-level symbolic search and low-level neural sampling in bilevel planning, IVNTR interleaves symbolic effect learning and neural classifier learning in an iterative process. In each iteration, the symbolic learning proposes a candidate predicate effect across different actions, which provides labels for neural learning on transition pairs. Once the neural classifier converges, its validation loss guides the symbolic learning to propose the next candidate that could minimize the loss in the new iteration. This iterative bilevel learning ultimately yields a compact set of neural predicates, which are then selected to optimize the planning objective [17]. The final set of invented predicates seamlessly integrates into operator and sampler learning frameworks [32, 17], ultimately forming a fully functional bilevel planner.
@@
To evaluate the effectiveness of IVNTR, we conduct extensive experiments across six diverse robot planning domains. These domains feature a wide range of low-level state representations, from SE(2) and SE(3) poses to high-dimensional point clouds. Furthermore, as shown in Figure LABEL:fig:teaser, by leveraging relational predicates and AI planning, IVNTR zero-shot generalizes to tasks with unseen entity compositions. Finally, we deploy IVNTR on a quadruped mobile manipulator (Boston Dynamics Spot) for two long-horizon mobile manipulation tasks. The learned predicates successfully abstract complex continuous states into representations compatible with the AI planner, while also providing actionable guidance for the samplers. We believe IVNTR represents a pivotal step towards learning high-level abstractions from sophisticated low-level states.
@@ II Problem Formulation
A solution to a task is a plan $\pi=[\underline{\mathtt{C}}_{1},\underline{\mathtt{C}}_{2},\cdots,\underline{\mathtt{C}}_{H}]$, that is, a sequence of $H$ ground actions such that successive application of the transition model $\mathbf{x}_{i}=f\left(\mathbf{x}_{i-1},\underline{\mathtt{C}}_{i}\right)$ on each $\underline{\mathtt{C}}_{i}\in\pi$, starting from $\mathbf{x}_{0}$, results in a final state $\mathbf{x}_{H}$ where $g$ holds. For instance, the plan depicted in Figure LABEL:fig:teaser bottom right finally leads to the state where $\texttt{In}(\texttt{t}_{1},\texttt{t}_{2})$ holds. In sum, to generate the plan $\pi$ for a task, in each state, the robot needs to predict: (1) the action class $\mathtt{C}\in\mathcal{C}$, (2) the objects as the discrete parameters, and (3) the continuous parameter $\omega\in\Omega$.
During training, the robot has access to an offline demonstration dataset $\mathcal{D}=\{({T}_{i},\pi_{i})\}_{i=1}^{B}$, which consists of $B$ task and solution pairs sampled from the task distribution $\mathcal{T}^{\mathrm{train}}$. Note that since the transition function $f$ is known and deterministic, we can also recover the intermediate states from $\mathbf{x}_{0}$. During test time, the robot is required to solve held-out tasks sampled from a different test distribution $\mathcal{T}^{\mathrm{test}}$. In practice, for the sake of evaluating generalization, test tasks typically contain new and more objects than training tasks. For example, as shown in Figure 1, all training tasks only have $1$ platform, but during test, there are $2$ platforms. The difference in object compositions could result in different lengths of plans with a different action order, requiring the method to generalize by understanding the implicit concepts present in the training demonstrations.
@@ III Bilevel Planning
Figure 2: (a) Overview of IVNTR during training. Given transition pairs in the continuous space, IVNTR invents neural predicates with different arguments parallelly, resulting in a candidate set. A subset that minimizes the planning objective $J(\cdot)$ is selected from the candidates, which serves as the final $\Psi_{\mathrm{dyn}}$. With the complete predicate set, sampler and operator learning can be achieved. (b) Bilevel planning with operators and samplers during test. Compositional ground predicates serve as inputs to the AI planner and provide guidance to the samplers.
@@
The operator for a parametrized controller $\mathtt{C}$ is a tuple, $\mathtt{Op}^{\mathtt{C}}=\langle\mathtt{Var},\mathtt{Pre},\mathtt{Eff}^{+},\mathtt{Eff}^{-}\rangle$, where $\mathtt{Var}$ is a tuple of object placeholders matching the type signature of $\mathtt{C}$, and $\mathtt{Pre},\,\mathtt{Eff}^{+},\,\mathtt{Eff}^{-}\subseteq\Psi$, respectively preconditions, add effects, and delete effects, are each a set of lifted predicates defined with variables in $\mathtt{Var}$. $\Psi$ is a predicate set.
@@
Given a task $T=\langle\mathcal{O},\mathbf{x}_{0},g\rangle$, bilevel planning (Figure 2b) starts by using predicates to generate an abstract state consisting of all ground predicates with objects $\mathcal{O}$ whose classifiers evaluate to True in $\mathbf{x}_{0}$. The initial ground predicates, together with the operator set and goal $g$, can then be input to an AI planner [34] to generate a plan skeleton, $\bar{\pi}$ with partially grounded actions with unspecified continuous parameters, $\underline{\bar{\mathtt{C}}}$. To refine the actions in this skeleton $\underline{\bar{\mathtt{C}}}\in\bar{\pi}$ into fully grounded $\underline{{\mathtt{C}}}$ with the continuous parameters $\omega$, bilevel planning leverages samplers.
@@
Note that unlike the deterministic operators, samplers are usually stochastic and may fail in certain steps. Thus, bilevel planning alternates between the AI planner and samplers, using the predicate classifiers $\theta_{\Psi}$ as “guidance” in each step to compensate for potential sampling failures.
Assuming we have a complete predicate set, previous work has studied the problem of learning operators [32] and samplers [19, 16] from the demonstration dataset $\mathcal{D}$. Since the predicates, learned operators, and samplers are relational, they can be generally applied to held-out test tasks sampled from $\mathcal{T}^{\mathrm{test}}$. However, with an insufficient predicate set—for example, with only $\Psi_{\mathrm{G}}$ and $\Psi_{\mathrm{sta}}$—bilevel planning can be intractably slow [17]. We next introduce the IVNTR framework, which closes this gap by automatically inventing dynamic predicates for efficient bilevel planning.
@@ IV Methodology
The problem of inventing dynamic predicates $\Psi_{\mathrm{dyn}}$ can be decomposed into symbolic learning—how many predicates should be invented, and with what type signatures—and classifier learning, determining $\theta_{\psi}$ for each invented predicate $\psi\in\Psi_{\mathrm{dyn}}$. Previous approaches [21, 17] address these problems via a “define-then-select” two-stage pipeline. In the first stage, for each predicate candidate $\hat{\psi}$ with known variable types, its classifer is pre-defined via program synthesis [17] or pre-trained foundation models [21]. These candidates form a large predicate pool $\hat{\Psi}_{\mathrm{dyn}}$. In the second stage, to subselect predicates from the pool, each candidate predicate set $\tilde{\Psi}_{\mathrm{dyn}}\subseteq\hat{\Psi}_{\mathrm{dyn}}$ is scored with a function $J(\tilde{\Psi}_{\mathrm{dyn}})$ that measures both planning efficiency and effectiveness. A key limitation of this “define-then-select” pipeline is that the classifiers within $\hat{\Psi}_{\mathrm{dyn}}$ are restricted to a relatively simple set. Scaling to more general classifiers, e.g., neural networks, is nontrivial, since $J(\tilde{\Psi}_{\mathrm{dyn}})$ is highly non-differentiable. To address this, we propose IVNTR, a “learn-then-select” approach that leverages bilevel learning.
As depicted in Figure 2 (a), given the domain types $\Lambda$, IVNTR enumerates all possible typed variable compositions (with maximum input arity). Since the input features for the predicate classifier depend on its argument types, IVNTR invents predicates with different arguments group by group parallelly. For the invention of each group, IVNTR draws inspiration from bilevel planning itself, where planning alternates between the symbolic level and the low level. Similarly, IVNTR interleaves symbolic learning and neural learning, with each providing guidance for the other. Specifically, symbolic learning proposes effect vectors that represent the add and delete effects for candidate predicates across all operators. Neural learning uses these effect vectors to provide supervision for classifier learning. The validation loss in neural learning feeds back into symbolic learning, and the process repeats. In this section, we describe these steps in detail via the exemplar predicate group $\psi\in\Psi^{\mathtt{Var}}$, with the typed variables $\mathtt{Var}=\mathtt{(?r,?t)}$.